Consider the following algebraic version of the IS-LM model C 200+0,5 YD (C is consumption, YD is disposable income); G (public spending) 100; T (taxes) 100 1. 350 4000i +0,1 Y (I is investment, i is interest rate, Y is real income). Real demand for money: md-0,5 Y-7500, real money supply: (MVP-mf-250; (i) Write the equations that represent the IS and LM relations. [3pl (ii Find the equilibrium values of real GDP and interest rate. [3p] (i) Find the equilibrium values of consumption, investment, and private saving. [3p] (iv) Let suppose a negative shock on autonomous consumption, which decreases from 200 to 100. If the central bank wants to stabilize output at the current level, at which value should it set the new money supply? (hint: first, compute the target interest rate, then the money supply). [3p]

PLEASE ANSWER ALL THE POINTS AND WRITE CLEARLY AND JUSTIFY THANKS
Answer #1

C+ TG 200 0.5 о.ЧУ = 600-4000 l LM curve = 500 + ISO001 Soure 1000 đS000 1 seoo&00 ± Ο.5 С 1100-100 350-160ナ1,0 an daing100 100- 300 T-C G00 Ч 530 t o.sy-SO-yo001.-to.Ly y ; laso-100001 Temau xed at liodo